OpenStudy (anonymous):

The point (-4, -1) is translated (x + 5, y + 2). Where is the translated point located? I want to say (-1,1) but I'm not sure.

OpenStudy (campbell_st):

not quite x = -4 + 5 y = -1 + 2 hope it helps

OpenStudy (anonymous):

So (1,1)?

OpenStudy (campbell_st):

that's it

OpenStudy (anonymous):

Thanks